OpenStudy (anonymous):

How to solve: 18v divided by 3v^2 ---- ----- t^2 - 4t + 3 6t - 6

OpenStudy (anonymous):

\[\frac{ 18v }{ t^2 - 4t + 3 } \div \frac{ 3v ^2 }{ 6t - 6 }\]

OpenStudy (harsimran_hs4):

lets start step 1: factorize t^2 -4t +3 ans tell me what you get

OpenStudy (anonymous):

(t - 3) (t - 1)

OpenStudy (harsimran_hs4):

great... and you know (a/b) / (c/d) = (a/b) * (d/c) so can you cancel the common factors now and get to the answer?

OpenStudy (anonymous):

But there are no common factors out of \[\frac{ 18v }{ (t-3)(t - 1) } \times \frac{ 6t - 6 }{ 3v^2 }\]

OpenStudy (harsimran_hs4):

i can see that v can cancel out from 18v and 3v^2 and (t-1) also because 6t-6 is 6*(t-1)

OpenStudy (anonymous):

\[\frac{ 18v }{ (t-3)(t-1) }\div \frac{ 3v^2 }{ 6(t-1) }\]\[\frac{ 18v }{ (t-3)(t-1) }\times \frac{ 6(t-1) }{ 3v^2 }\]Cancels (t-1), 18 becomes 6 and 3 becomes 1, v cancels in the numerator and v2 becomes v\[\frac{ 6 }{ t-3}\times \frac{ 6 }{ v }\]\[\frac{ 36 }{ v(t-3) }\]\[\frac{ 36 }{ tv-3v }\]

OpenStudy (harsimran_hs4):

@haileemackk i wanted you to figure out the steps but Lynncake did all the steps so clear now?

OpenStudy (anonymous):

i think i understand now. thank you for your help